# Biorthogonal measures, polymer partition functions, and random matrices

### v2

@@ 5 Log Gamma polymer partition function and LUE with external source
$(1,1)$$(n,N)$
$\alpha_{1}$$\alpha_{2}$$\alpha_{3}$$\alpha_{4}$$\alpha_{5}$$\alpha_{6}$$\alpha_{7}$$\alpha_{8}$$\alpha_{9}$
$a_{1}$$a_{2}$$a_{3}$$a_{4}$$a_{5}$$a_{6}$
@@
As anticipated in the introduction, Remark 1.9, combining the Fredholm determinant identity (5.3) with Theorem 1.7, part 4, we can relate the partition function of the Log Gamma polymer to an explicit biorthogonal measure. In the following corollary and below, given $b\in\mathbb{C}$, $[b]_{k}$ will denote a $k$-dimensional vector with $b$ in each entry and $\vec{e}_{m}$ will denote the $N$-dimensional vector containing $1$ in position $m$ and 00 elsewhere.
@@
    In the case where the parameters $\vec{a}$ are all distinct, the biorthogonal measure is given explicitly, in terms of Meijer $G$-function [31], by
@@
Let $A$ be a deterministic positive-definite Hermitian $N\times N$ matrix with eigenvalues $a_{1},\ldots,a_{N}\in[0,1)$, and consider the probability distribution
$\frac{1}{Z_{N}^{\prime}}(\det M)^{\nu}{\rm e}^{-{\rm Tr}((I-A)M)}{\rm d}M,\qquad\nu>0,\quad{\rm d}M=\prod_{j=1}^{N}{\rm d}M_{jj}\ \prod_{1\leq j<k\leq N}{\rm d}\mathrm{Re}\,M_{jk}\,{\rm d}\mathrm{Im}\,M_{jk},$
on the space of $N\times N$ positive-definite Hermitian matrices. This model is known as the chiral Unitary Ensemble or LUE with external source and is connected to sample covariance matrices, see e.g. [4, 29, 32, 33]. For $A=0$, this is the classical LUE or Wishart-Laguerre ensemble.
@@
$\frac{1}{Z_{N}}\Delta(\vec{x})\ \det\left(x_{j}^{\nu}{\rm e}^{-(1-a_{k})x_{j}}\right)_{j,k=1}^{N}\prod_{j=1}^{N}{\rm d}x_{j},\qquad x_{1},\ldots,x_{N}>0.$ (5.11)
@@
$L_{N}^{\rm LUE+}(x,x^{\prime};\vec{a},\nu):=\frac{{\rm e}^{x^{\prime}-x}}{(2\pi\mathrm{i})^{2}}\int_{\gamma_{1}}{\rm d}u\int_{\gamma_{2}}{\rm d}v\left(\frac{u}{v}\right)^{N+\nu}\prod_{j=1}^{N}\frac{v-1+a_{j}}{u-1+a_{j}}\frac{{\rm e}^{-ux^{\prime}+vx}}{v-u},$ (5.12)
where $\gamma_{1}$ encircles $1-a_{1},\ldots,1-a_{N}$ anti-clockwise, $\gamma_{2}$ encircles 00 anti-clockwise, and they don’t intersect. Changing $u\mapsto 1-u$, $v\mapsto 1-v$, this is equal to
$L_{N}^{\rm LUE+}(x,x^{\prime};\vec{a},\nu):=\frac{1}{(2\pi\mathrm{i})^{2}}\int_{\Sigma_{N}}{\rm d}u\int_{\ell_{N}}{\rm d}v\left(\frac{u-1}{v-1}\right)^{N+\nu}\prod_{j=1}^{N}\frac{v-a_{j}}{u-a_{j}}\frac{{\rm e}^{-vx+ux^{\prime}}}{v-u},$ (5.13)
where $\Sigma_{N}$ is a positively oriented circle around $a_{1},\ldots,a_{N}$ with radius $<1$, and $\ell_{N}$ a negatively oriented circle around $1$, which can moreover be deformed (after changing the orientation) to a vertical line lying between $\Sigma_{N}$ and $1$ if $\nu>0$. This kernel is of the form (1.8) with
$W_{N}(z)\equiv W_{N}^{\rm LUE+}(z;\vec{a},\nu):=\frac{\prod_{j=1}^{N}(z-a_{j})}{(z-1)^{N+\nu}}.$ (5.14)
@@
and re-scaled $a$-parameters $Ta_{1},\ldots,Ta_{N}$ with $10\leq a_{1},\ldots,a_{N}<$ 1. We then have
$W^{{\rm Log}\Gamma}_{n,N}(T\zeta;[T]_{n},T\vec{a})$ $=\frac{\Gamma(T(1-\zeta))^{n}}{\prod_{m=1}^{N}\Gamma(T(\zeta-a_{m}))}$
$=T^{N-n}\frac{\prod_{m=1}^{N}(\zeta-a_{m})}{(1-\zeta)^{n}}\frac{\Gamma(1+T(1-\zeta))^{n}}{\prod_{m=1}^{N}\Gamma(1+T(\zeta-a_{m}))}.$
@@
$\lim_{T\to 0}\frac{W^{{\rm Log}\Gamma}_{n,N}(Tv;[T]_{n},T\vec{a})}{W^{{\rm Log}\Gamma}_{n,N}(Tu;[T]_{n},T{\vec{a}})}=\frac{W_{N}^{\rm LUE+}(v;{\vec{a}},n-N)}{W_{N}^{\rm LUE+}(u;{\vec{a}},n-N)},\qquad\mbox{for $v\in\ell_{N}$, $u\in\Sigma_{N}$}.$ (5.15)
@@
$\lim_{T\to 0}\frac{1}{T}L_{n,N}^{{\rm Log}\Gamma}\left(\frac{x}{T},\frac{x^{\prime}}{T};[T]_{n},T{\vec{a}}\right)=L_{N}^{\rm LUE+}\left(x,x^{\prime};{\vec{a}},\nu=n-N\right),\qquad x,x^{\prime}\in\mathbb{R},$ (5.16)
and there exists $C_{n,N}(\vec{a})$ such that for $\epsilon,T>0$ sufficiently small,
$\left|\frac{1}{T}L_{n,N}^{{\rm Log}\Gamma}\left(\frac{x}{T},\frac{x^{\prime}}{T};[T]_{n},T{\vec{a}}\right)\right|\leq C_{n,N}(\vec{a})h(x)\tilde{h}(x^{\prime}),$ (5.17)
@@
$h(x)={\rm e}^{-(a_{\max}+2\epsilon)x},\qquad\tilde{h}(x^{\prime})=\begin{cases}{\rm e}^{(a_{\min}-\epsilon)x^{\prime}},&x^{\prime}<0\\ {\rm e}^{(a_{\max}+\epsilon)x^{\prime}},&x^{\prime}\geq 0\end{cases}.$ (5.18)
@@
$\frac{1}{T}L_{n,N}^{{\rm Log}\Gamma}\left(\frac{x}{T},\frac{x^{\prime}}{T};[T]_{n},T{\vec{a}}\right)=\frac{1}{(2\pi\mathrm{i})^{2}}\int_{\Sigma_{N}}{\rm d}s\int_{\ell_{N}}{\rm d}w\frac{W_{n,N}^{{\rm Log}\Gamma}(Tw;[T]_{n},T\vec{a})}{W_{n,N}^{{\rm Log}\Gamma}(Ts;[T]_{n},T\vec{a})}\frac{{\rm e}^{-wx+sx^{\prime}}}{w-s}$
$\quad=\frac{1}{(2\pi\mathrm{i})^{2}}\int_{\Sigma_{N}}{\rm d}s\int_{\ell_{N}}{\rm d}w\left(\frac{s-1}{w-1}\right)^{n}\frac{\Gamma(1+T(1-w))^{n}}{\Gamma(1+T(1-s))^{n}}\prod_{j=1}^{N}\frac{w-a_{j}}{s-a_{j}}\frac{\Gamma(1+T(s-a_{j}))}{\Gamma(1+T(w-a_{j}))}\frac{{\rm e}^{-wx+sx^{\prime}}}{w-s},$
where $\Sigma_{N}$ is a contour surrounding $a_{1},\ldots,a_{N}$ and no other zeros of $W_{n,N}^{{\rm Log}\Gamma}(Tu)$, and $\ell_{N}$ lies at the right of $\Sigma_{N}$ and at the left of $1$, and where we used $\Gamma(z+1)=z\Gamma(z)$ to pass to the second line. Hence,
$\frac{1}{T}L_{n,N}^{{\rm Log}\Gamma}\left(\frac{x}{T},\frac{x^{\prime}}{T};[T]_{n},T{\vec{a}}\right)$
$\quad=\frac{1}{(2\pi\mathrm{i})^{2}}\int_{\Sigma_{N}}{\rm d}s\int_{\ell_{N}}{\rm d}w\frac{W_{N}^{\rm LUE+}(w;\vec{a},n-N)}{W_{N}^{\rm LUE+}(s;\vec{a},n-N)}\left(\frac{\Gamma(1+T(1-w))^{n}}{\Gamma(1+T(1-s))^{n}}\prod_{j=1}^{N}\frac{\Gamma(1+T(s-a_{j}))}{\Gamma(1+T(w-a_{j}))}\right)\frac{{\rm e}^{-wx+sx^{\prime}}}{w-s}.$
@@
To obtain the bound (5.17), we take $\ell_{N}=a_{\max}+2\epsilon+\mathrm{i}\mathbb{R}$, and $\Sigma_{N}$ within the strip $a_{\min}-\epsilon\leq\mathrm{Re}\,s\leq a_{\max}+\epsilon$, with $\epsilon>0$ sufficiently small so that $a_{\max}+2\epsilon<1$. For $T>0$ sufficiently small, it then follows that
$\left|\frac{1}{T}L_{n,N}^{{\rm Log}\Gamma}\left(\frac{x}{T},\frac{x^{\prime}}{T};[T]_{n},T{\vec{a}}\right)\right|$ $\leq 2\frac{h(x)\tilde{h}(x^{\prime})}{(2\pi)^{2}}\int_{\Sigma_{N}}|{\rm d}s|\int_{\ell_{N}}|{\rm d}w|\left|\frac{W_{N}^{\rm LUE+}(w;\vec{a},n-N)}{W_{N}^{\rm LUE+}(s;\vec{a},n-N)}\right|\frac{1}{|w-s|}$
$\leq C_{n,N}({\vec{a}})h(x)\widetilde{h}(x^{\prime}).$
The latter inequality holds only if $n>N$, since $\left|\frac{W_{N}^{\rm LUE+}(w;\vec{a},n-N)}{W_{N}^{\rm LUE+}(s;\vec{a},n-N)}\right|\frac{1}{|w-s|}$ is not integrable for $n=N$. One can however refine the above estimates by taking into account the decay of $\frac{\Gamma(1+T(1-w))^{n}}{\Gamma(1+T(1-s))^{n}}\prod_{j=1}^{N}\frac{\Gamma(1+T(s-a_{j}))}{\Gamma(1+T(w-a_{j}))}$ as $w\to\infty$ on $\ell_{N}$, such that the same inequality remains to hold. ∎
@@
$\mathbb{E}\left[{\rm e}^{-{\rm e}^{t/T}Z_{n,N}^{{\rm Log}\Gamma}([T]_{n},T{\vec{a}})}\right]\longrightarrow\mathbb{P}_{\mathrm{LUE+}}\left(\max\{x_{1},\ldots,x_{N}\}\leq-t;\vec{a},\nu=N-n\right),$ (5.19)
@@
$\det\left(\frac{\sigma_{t/T}(x_{j}/T)}{T}L_{n,N}^{{\rm Log}\Gamma}\left(\frac{x_{j}}{T},\frac{x_{k}}{T};[T]_{n},T{\vec{a}}\right)\right)_{j,k=1}^{N}\to\det\left(1_{(-t,+\infty)}(x)L_{N}^{\rm LUE+}\left(x_{j},x_{k};{\vec{a}},n-N\right)\right)_{j,k=1}^{N}$
@@
$\left|\det\left(\frac{\sigma_{t/T}(x_{j}/T)}{T}L_{n,N}^{{\rm Log}\Gamma}\left(\frac{x_{j}}{T},\frac{x_{k}}{T};[T]_{n},T{\vec{a}}\right)\right)_{j,k=1}^{N}\right|\leq C_{n,N}^{N}\prod_{j=1}^{N}h(x_{j})\tilde{h}(x_{j})\sigma_{t/T}(x_{j}/T)\ |\det O(1)|,$

### v3

@@ 5 Log Gamma polymer partition function and LUE with external source
$(1,1)$$(n,N)$$\alpha_{1}$$\alpha_{2}$$\alpha_{3}$$\alpha_{4}$$\alpha_{5}$$\alpha_{6}$$\alpha_{7}$$\alpha_{8}$$\alpha_{9}$$a_{1}$$a_{2}$$a_{3}$$a_{4}$$a_{5}$$a_{6}$
@@
As anticipated in the introduction, Remark 1.9, combining the Fredholm determinant identity (5.3) with Theorem 1.7, part 4, we can relate the partition function of the Log Gamma polymer to an explicit biorthogonal measure. In the following corollary and below, given $b\in\mathbb{C}$, $[b]_{k}$ will denote a $k$-dimensional vector with $b$ in each entry and $\vec{e}_{m}$ will denote the $N$-dimensional vector containing $1$ in position $m$ and 00 elsewhere. We also recall the definition of Meijer $G$-function (see for instance [31])
$G^{\,*,m}_{n,p}\lparen\begin{smallmatrix}q\\ a_{1},\ldots,a_{p}\end{smallmatrix}|\,b_{1},\ldots,b_{q}\rparen{z}:=\frac{1}{2\pi\mathrm{i}}\int_{L}\frac{\prod_{\ell=1}^{m}\Gamma(b_{\ell}-v)\prod_{\ell=1}^{n}\Gamma(1-a_{\ell}+v)}{\prod_{\ell=m}^{q-1}\Gamma(1-b_{\ell+1}+v)\prod_{\ell=n}^{p-1}\Gamma(a_{\ell+1}-v)}z^{v}{\rm d}v.$
The choice of the contour $L$ depends on the positions of the parameters $\{a_{i},\,i=1,\ldots,p\}$ and $\{b_{j},\,j=1,\ldots,q\}$. For our specific cases below, we can choose it as the vertical line $\ell_{N}$.
@@
    In the case where the parameters $\vec{a}$ are all distinct, the biorthogonal measure is given explicitly, in terms of Meijer $G$-function, by
@@
Remark 5.3.
It is easy to see, using the definition of Meijer $G$-functions, that when $W_{N}$ is given by (5.4), the associated functions $\{\psi_{i}\}_{i=1}^{N}$ and $\{\phi_{i}\}_{i=1}^{N}$, defined by (1.12) and (1.17), are real valued:
$\psi_{i}({\rm e}^{\bar{x}})=\overline{\psi_{i}({\rm e}^{x})},\quad\phi_{i}({\rm e}^{\bar{x}})=\overline{\phi_{i}({\rm e}^{x})},\quad i=1,\ldots,N.$
Consequently, $d\mu_{n,N}^{{\rm Log}\Gamma}$ is a signed real-valued measure.
@@
Let $B$ be a deterministic positive-definite Hermitian $N\times N$ matrix with eigenvalues $b_{1},\ldots,b_{N}\in[0,1)$, and consider the probability distribution
$\frac{1}{Z_{N}^{\prime}}(\det M)^{\nu}{\rm e}^{-{\rm Tr}((I-B)M)}{\rm d}M,\qquad\nu>0,\quad{\rm d}M=\prod_{j=1}^{N}{\rm d}M_{jj}\ \prod_{1\leq j<k\leq N}{\rm d}\mathrm{Re}\,M_{jk}\,{\rm d}\mathrm{Im}\,M_{jk},$
on the space of $N\times N$ positive-definite Hermitian matrices. This model is known as the chiral Unitary Ensemble or LUE with external source and is connected to sample covariance matrices, see e.g. [4, 29, 32, 33]. For $B=0$, this is the classical LUE or Wishart-Laguerre ensemble.
@@
$\frac{1}{Z_{N}}\Delta(\vec{x})\ \det\left(x_{j}^{\nu}{\rm e}^{-(1-b_{k})x_{j}}\right)_{j,k=1}^{N}\prod_{j=1}^{N}{\rm d}x_{j},\qquad x_{1},\ldots,x_{N}>0.$ (5.11)
@@
$L_{N}^{\rm LUE+}(x,x^{\prime};\vec{b},\nu):=\frac{{\rm e}^{x^{\prime}-x}}{(2\pi\mathrm{i})^{2}}\int_{\gamma_{1}}{\rm d}u\int_{\gamma_{2}}{\rm d}v\left(\frac{u}{v}\right)^{N+\nu}\prod_{j=1}^{N}\frac{v-1+b_{j}}{u-1+b_{j}}\frac{{\rm e}^{-ux^{\prime}+vx}}{v-u},\qquad x,x^{\prime}>0,$ (5.12)
where $\gamma_{1}$ encircles $1-b_{1},\ldots,1-b_{N}$ anti-clockwise, $\gamma_{2}$ encircles 00 anti-clockwise, and they don’t intersect. Changing $u\mapsto 1-u$, $v\mapsto 1-v$, this is equal to
$L_{N}^{\rm LUE+}(x,x^{\prime};\vec{b},\nu):=\frac{1}{(2\pi\mathrm{i})^{2}}\int_{\Sigma_{N}}{\rm d}u\int_{\ell_{N}}{\rm d}v\left(\frac{u-1}{v-1}\right)^{N+\nu}\prod_{j=1}^{N}\frac{v-b_{j}}{u-b_{j}}\frac{{\rm e}^{-vx+ux^{\prime}}}{v-u},$ (5.13)
where $\Sigma_{N}$ is a positively oriented circle around $b_{1},\ldots,b_{N}$ with radius $<1$, and $\ell_{N}$ a negatively oriented circle around $1$, which can moreover be deformed (after changing the orientation) to a vertical line lying between $\Sigma_{N}$ and $1$ if $\nu>0$. This kernel is of the form (1.8) with
$W_{N}(z)\equiv W_{N}^{\rm LUE+}(z;\vec{b},\nu):=\frac{\prod_{j=1}^{N}(z-b_{j})}{(z-1)^{N+\nu}}.$ (5.14)
@@
and re-scaled $a$-parameters $a_{j}=Tb_{j},\quad j=1\ldots,N$ with $10\leq b_{1},\ldots,b_{N}<$ 1. We then have
$W^{{\rm Log}\Gamma}_{n,N}(T\zeta;[T]_{n},T\vec{b})$ $=\frac{\Gamma(T(1-\zeta))^{n}}{\prod_{m=1}^{N}\Gamma(T(\zeta-b_{m}))}$
$=T^{N-n}\frac{\prod_{m=1}^{N}(\zeta-b_{m})}{(1-\zeta)^{n}}\frac{\Gamma(1+T(1-\zeta))^{n}}{\prod_{m=1}^{N}\Gamma(1+T(\zeta-b_{m}))}.$
@@
$\lim_{T\to 0}\frac{W^{{\rm Log}\Gamma}_{n,N}(Tv;[T]_{n},T\vec{b})}{W^{{\rm Log}\Gamma}_{n,N}(Tu;[T]_{n},T\vec{b})}=\frac{W_{N}^{\rm LUE+}(v;\vec{b},n-N)}{W_{N}^{\rm LUE+}(u;\vec{b},n-N)},\qquad\mbox{for $v\in\ell_{N}$, $u\in\Sigma_{N}$}.$ (5.15)
@@
$\lim_{T\to 0}\frac{1}{T}L_{n,N}^{{\rm Log}\Gamma}\left(\frac{x}{T},\frac{x^{\prime}}{T};[T]_{n},T\vec{b}\right)=L_{N}^{\rm LUE+}\left(x,x^{\prime};\vec{b},\nu=n-N\right),\qquad x,x^{\prime}\in\mathbb{R},$ (5.16)
and there exists $C_{n,N}(\vec{b})$ such that for $\epsilon,T>0$ sufficiently small,
$\left|\frac{1}{T}L_{n,N}^{{\rm Log}\Gamma}\left(\frac{x}{T},\frac{x^{\prime}}{T};[T]_{n},T\vec{b}\right)\right|\leq C_{n,N}(\vec{b})h(x)\tilde{h}(x^{\prime}),$ (5.17)
@@
$h(x)={\rm e}^{-(b_{\max}+2\epsilon)x},\qquad\tilde{h}(x^{\prime})=\begin{cases}{\rm e}^{(b_{\min}-\epsilon)x^{\prime}},&x^{\prime}<0\\ {\rm e}^{(b_{\max}+\epsilon)x^{\prime}},&x^{\prime}\geq 0\end{cases}.$ (5.18)
@@
$\frac{1}{T}L_{n,N}^{{\rm Log}\Gamma}\left(\frac{x}{T},\frac{x^{\prime}}{T};[T]_{n},T\vec{b}\right)=\frac{1}{(2\pi\mathrm{i})^{2}}\int_{\Sigma_{N}}{\rm d}s\int_{\ell_{N}}{\rm d}w\frac{W_{n,N}^{{\rm Log}\Gamma}(Tw;[T]_{n},T\vec{b})}{W_{n,N}^{{\rm Log}\Gamma}(Ts;[T]_{n},T\vec{b})}\frac{{\rm e}^{-wx+sx^{\prime}}}{w-s}$
$\quad=\frac{1}{(2\pi\mathrm{i})^{2}}\int_{\Sigma_{N}}{\rm d}s\int_{\ell_{N}}{\rm d}w\left(\frac{s-1}{w-1}\right)^{n}\frac{\Gamma(1+T(1-w))^{n}}{\Gamma(1+T(1-s))^{n}}\prod_{j=1}^{N}\frac{w-b_{j}}{s-b_{j}}\frac{\Gamma(1+T(s-b_{j}))}{\Gamma(1+T(w-b_{j}))}\frac{{\rm e}^{-wx+sx^{\prime}}}{w-s},$
where $\Sigma_{N}$ is a contour surrounding $b_{1},\ldots,b_{N}$ and no other zeros of $W_{n,N}^{{\rm Log}\Gamma}(Tu)$, and $\ell_{N}$ lies at the right of $\Sigma_{N}$ and at the left of $1$, and where we used $\Gamma(z+1)=z\Gamma(z)$ to pass to the second line. Hence,
$\frac{1}{T}L_{n,N}^{{\rm Log}\Gamma}\left(\frac{x}{T},\frac{x^{\prime}}{T};[T]_{n},T\vec{b}\right)$
$\quad=\frac{1}{(2\pi\mathrm{i})^{2}}\int_{\Sigma_{N}}{\rm d}s\int_{\ell_{N}}{\rm d}w\frac{W_{N}^{\rm LUE+}(w;\vec{b},n-N)}{W_{N}^{\rm LUE+}(s;\vec{b},n-N)}\left(\frac{\Gamma(1+T(1-w))^{n}}{\Gamma(1+T(1-s))^{n}}\prod_{j=1}^{N}\frac{\Gamma(1+T(s-b_{j}))}{\Gamma(1+T(w-b_{j}))}\right)\frac{{\rm e}^{-wx+sx^{\prime}}}{w-s}.$
@@
To obtain the bound (5.17), we take $\ell_{N}=b_{\max}+2\epsilon+\mathrm{i}\mathbb{R}$, and $\Sigma_{N}$ within the strip $b_{\min}-\epsilon\leq\mathrm{Re}\,s\leq b_{\max}+\epsilon$, with $\epsilon>0$ sufficiently small so that $b_{\max}+2\epsilon<1$. For $T>0$ sufficiently small, it then follows that
$\left|\frac{1}{T}L_{n,N}^{{\rm Log}\Gamma}\left(\frac{x}{T},\frac{x^{\prime}}{T};[T]_{n},T\vec{b}\right)\right|$ $\leq 2\frac{h(x)\tilde{h}(x^{\prime})}{(2\pi)^{2}}\int_{\Sigma_{N}}|{\rm d}s|\int_{\ell_{N}}|{\rm d}w|\left|\frac{W_{N}^{\rm LUE+}(w;\vec{b},n-N)}{W_{N}^{\rm LUE+}(s;\vec{b},n-N)}\right|\frac{1}{|w-s|}$
$\leq C_{n,N}(\vec{b})h(x)\widetilde{h}(x^{\prime}).$
The latter inequality holds only if $n>N$, since $\left|\frac{W_{N}^{\rm LUE+}(w;\vec{b},n-N)}{W_{N}^{\rm LUE+}(s;\vec{b},n-N)}\right|\frac{1}{|w-s|}$ is not integrable for $n=N$. One can however refine the above estimates by taking into account the decay of $\frac{\Gamma(1+T(1-w))^{n}}{\Gamma(1+T(1-s))^{n}}\prod_{j=1}^{N}\frac{\Gamma(1+T(s-b_{j}))}{\Gamma(1+T(w-b_{j}))}$ as $w\to\infty$ on $\ell_{N}$, such that the same inequality remains to hold. ∎
@@
$\mathbb{E}\left[{\rm e}^{-{\rm e}^{t/T}Z_{n,N}^{{\rm Log}\Gamma}([T]_{n},T\vec{b})}\right]\longrightarrow\mathbb{P}_{\mathrm{LUE+}}\left(\max\{x_{1},\ldots,x_{N}\}\leq-t;\vec{b},\nu={n-N}\right),$ (5.19)
@@
$\det\left(\frac{\sigma_{t/T}(x_{j}/T)}{T}L_{n,N}^{{\rm Log}\Gamma}\left(\frac{x_{j}}{T},\frac{x_{k}}{T};[T]_{n},T\vec{b}\right)\right)_{j,k=1}^{N}\to\det\left(1_{(-t,+\infty)}(x)L_{N}^{\rm LUE+}\left(x_{j},x_{k};\vec{b},n-N\right)\right)_{j,k=1}^{N}$
@@
$\left|\det\left(\frac{\sigma_{t/T}(x_{j}/T)}{T}L_{n,N}^{{\rm Log}\Gamma}\left(\frac{x_{j}}{T},\frac{x_{k}}{T};[T]_{n},T\vec{b}\right)\right)_{j,k=1}^{N}\right|\leq C_{n,N}^{N}\prod_{j=1}^{N}h(x_{j})\tilde{h}(x_{j})\sigma_{t/T}(x_{j}/T)\ |\det O(1)|,$